Find using distributive property 5437×1001 .

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
As we know that in the above question we have to find the product of numbers but by using the distributive property. We can say that the product is the result of one or more multiplications. We know that there are three properties of numbers that are very commonly used. They are Commutative property, Associative property and Distributive property. In this question we have to use the last one. Complete step-by-step answer: Given numbers in the question are 5437×1001 . Since distributive property means to divide the given operations on the numbers so that the equation becomes easier to solve. If there is a(b+c) , then we apply distributive property and write it as a×b+a×c=ab+ac . In this question we will break 1001 as (1000+1) . So the given term can be written as 5437(1000+1) . By applying the above formula we can solve: 5437(1000+1)=5437×1000+5437×1 . On further solving we have, 5437000+5437=5442437 . Hence the required value is 5442437 . So, the correct answer is “ 5442437 .”. Note: We should note that distributive property is also known as the law of multiplication. Distribute means the name itself implies that to divide something. It also meant “ multiplication distributed over addition.” This property can be applied over addition, subtraction, multiplication and Division.